hey,can you help me solve this polynomial using factoring.... i have tried numerous times and it doesnt work out right.....
6Xsquared+2XY+Ysquared+3XY

Answer by josmiceli(6786)

6Xsquared+2XY+Ysquared+3XY
or
6x^2 +2xy +y^2 + 3xy
combine the xy terms
6x^2 + 5xy + y^2
(2x + y)(3x + y)
pick some values for x and y to check answer
x = 5
y = -3
6(5)^2 + 5(5)(-3) + (-3)^2 =
150 + (-75) + 9 = 84
now check the factors
(2(5) + (-3))*(3(5) + (-3)) =
(10 + (-3))*(15 + (-3))
(7)*(12) = 84
OK